OpenStudy (anonymous):

**MULTIPLE CHOICE** The measures of the opposite angles of a parallelogram are represented by 2x+34 and 3x-12. Find the value of x. (1) 22 (2) 46 (3) 78 (4) 126

OpenStudy (anonymous):



OpenStudy (anonymous):

78

OpenStudy (anonymous):



OpenStudy (radar):

What do you know about opposite angles of a parallelogram?

OpenStudy (radar):

Earth to @CrackedCode ?

OpenStudy (radar):

Amy got it, they are indeed equal. That indeed is the key. Use algebra and set up the equality. 2x + 34 = 3x -12 Solve for x.

OpenStudy (radar):

34 + 12 = 3x - 2x 46 = 1x x=46